# Recurrence in a periodically driven and weakly damped Fermi-Pasta-Ulam-Tsingou chain

### v1

@@ Abstract
We report numerical evidence of Fermi-Pasta-Ulam-Tsingou (FPUT)–like recurrence in weakly damped, periodically driven $\alpha$-FPUT chains. In narrow regions of driving amplitude and damping, energy is quasi-periodically exchanged among a few low-frequency modes over long timescales. Unlike discrete time crystals, the recurrence period is not an integer multiple of the driving period and does not correspond to spontaneous symmetry breaking, yet it may suggests a generalized, relaxed form of time-crystalline-like order. The maximum damping allowing recurrence decreases rapidly with chain length, suggesting that in the thermodynamic limit such behavior is unlikely to persist. These results reveal a new type of coherent nonlinear dynamics in driven, open multimode systems and provide guidance for experimentally realizing long-lived quasi-periodic states.
@@ I Introduction
In this Letter, we present numerical evidence that a recurrence-like phenomenon can emerge in a periodically driven, weakly damped Fermi-Pasta-Ulam-Tsingou (FPUT) chain. Specifically, near the first mode, we identify a narrow region in the parameter space of driving amplitude and damping where energy is periodically exchanged among a few modes over long timescales. The maximum damping that allows such recurrence decreases rapidly with system size–for example, it is approximately $4.3\times 10^{-3}$ for $N=8$, and about $5\times 10^{-5}$ for $N=32$. These results suggest that in the thermodynamic limit, such recurrence behavior is unlikely to persist. We interpret this phenomenon as a new type of coherent nonlinear dynamics enabled by a delicate balance between weak driving, weak damping, and the system’s modal structure.
The paper is organized as follows. Section II introduces the model and the methodology used in our numerical simulations. Section III presents the numerical results, including representative recurrence phenomena and the dependence of the system’s response on driving frequency, driving amplitude, damping, and chain length. Finally, Sec. IV discusses the connection between the observed recurrence behavior and the recently explored concept of time crystals.
@@ II Model and methodology for our numerical simulations
$\ddot{q}_{n}=$ $(q_{n+1}-2q_{n}+q_{n-1})+\alpha\left[(q_{n+1}-q_{n})^{2}-(q_{n}-q_{n-1})^{2}\right]$ (1)
$-\eta\dot{q}_{n}+F\cos(\Omega t),\quad\quad(n=1,2,.,N),$
@@
We integrate Eq. (1) using a Runge-Kutta algorithm (verified with both MATLAB’s ode45 and Python’s scipy.integrate.solve_ivp, yielding consistent results), with initial conditions $q_{n}=0$ and $\dot{q}_{n}=0$. In the original FPUT problem, Runge-Kutta methods are generally not used because they are not symplectic and therefore do not guarantee long-term energy conservation. In our driven–damped model, however, symplectic algorithm is not required. To minimize the influence of numerical errors, we set the solver’s error tolerances smaller than the default values (e.g., $\text{relative error tolerance}=10^{-8}$, $\text{absolute error tolerance}=10^{-12}$).
we have monitored the hamiltonian of the undriven, undamped $\alpha$-FPUT as the ”total energy”
@@
$\begin{bmatrix}q_{n}\\ p_{n}\end{bmatrix}=\sqrt{\frac{2}{N+1}}\sum_{k=1}^{\mathrm{N}}\begin{bmatrix}\mathrm{Q}_{k}\\ \mathrm{P}_{k}\end{bmatrix}\sin\left(\frac{nk\pi}{\mathrm{N}+1}\right).$ (5)
Figure 2: Representative recurrence phenomenon. (a–c) Simulation results for the $N=8$ chain with $\eta=2.5\times 10^{-3}$, driving amplitude $F=5\times 10^{-3}$, and driving frequency $\omega=0.347$ (close to the first mode $\omega_{1}=0.3473$). The horizontal axis represents time in units of the driving period $T_{0}$. (d) Fast Fourier transform (FFT) of the steady-state response shown in panel (c).
@@ III Numerical results
Figure 3: (a) Total energy spectra of the $N=8$ chain as a function of the driving frequency near the first mode $\omega_{1}=0.3473$, for different driving amplitudes. (b) The same data for $F=5\times 10^{-3}$ shown separately, with a linear scale on the vertical axis. (c) Enlarged view of the frequency range where the recurrence phenomena occur, obtained from a finer frequency scan of the section shown in (b). (d–f) Time-domain simulations for three representative driving frequencies.
@@ III.1.2 Energy spectra versus driving frequency
Figure 3(a) shows the total energy response as a function of the driving frequency near the first mode, $\omega_{1}=0.3473$, for different driving amplitudes. Specifically, for each set of control parameters, Eq. 1 was integrated for $N$ driving periods $T_{0}$ (typically $N=2000$) until transients had decayed. The total energy was then stroboscopically sampled with period $T_{0}$ over the following $500\,T_{0}$, and these data points are plotted in Fig. 3(a). To maintain consistency, the first sampled point for each frequency scan was taken at the maximum total energy within one $T_{0}$ period.
As the driving amplitude increases, nonlinear effects become more pronounced, giving rise to secondary resonances near the primary peaks, accompanied by jump phenomena in their vicinity. The recurrence phenomena occur in the vicinity of these jumps, which is more clearly seen in Fig. 3(b) and (c). On both sides of the recurrence regions, the system still exhibits normal steady-state oscillations synchronized with the driving period (Figs. 3(d) and (f)).
@@
In addition to the uniform driving, we also examined two other types of excitation: a single-site driving $F_{n}=a\,\delta_{n,n_{0}}\sin(\Omega t)$ and a mode-matched driving $F_{n}=a\,\sin\!\left(\tfrac{nk\pi}{N+1}\right)\sin(\Omega t)$. Both cases produce similar recurrence behaviors, indicating that the recurrence is an intrinsic consequence of the system’s nonlinear mode coupling rather than a specific feature of the uniform excitation.
@@ III.1.3 Energy spectra versus driving amplitude
Figure 7: Recurrence regions in the driving amplitude–damping plane for chains of length $N=8$, 1616, and $32$, each driven at a frequency close to its first normal mode. Red filled circles and black filled squares indicate the boundary points of each region, and the connecting lines are drawn only to guide the eye. Note the logarithmic scale: the parameter space in which recurrence is observed becomes significantly narrower for $N=32$ compared to $N=8$.
@@ III.2 Recurrence regions versus chain Length
IV Discussion and conduction
We report numerical evidence of recurrence phenomena in $\alpha$-FPUT chains. Similar investigations were carried out for $\beta$-FPUT chains, but no analogous recurrence was found within the finite parameter space we explored ($\beta=8$). Moreover, since the driven FPUT problem is a prototypical multimode system with nonlinear mode coupling, deriving analytical approximations that describe the system’s behavior inside the recurrence regions remains challenging, providing a natural direction for future theoretical exploration.
@@ Acknowledgments
This research was supported by the Research Proiect Supported by Shanxi Scholarship Council of China (Grant No. 2023-033), and the Fundamental Research Program of Shanxi Province (Grant No. 202303021221071).

### v2

@@ Abstract
We report numerical evidence of Fermi-Pasta-Ulam-Tsingou (FPUT)–like recurrence in weakly damped, periodically driven $\alpha$-FPUT chains. In narrow regions of driving amplitude and damping, the steady-state energy is exchanged among a few low-frequency modes in a quasi-periodic (or highly regular, near-periodic) manner over long timescales. The maximum damping allowing recurrence decreases rapidly with chain length, suggesting that in the thermodynamic limit such behavior is unlikely to persist. Unlike discrete time crystals, the recurrence period is not an integer multiple of the driving period and does not correspond to spontaneous symmetry breaking. Nevertheless, these results reveal a new type of coherent nonlinear dynamics in driven, open multimode systems and provide guidance for experimentally realizing long-lived quasi-periodic states.
@@ I Introduction
In this Letter, we present numerical evidence that a recurrence-like phenomenon can emerge in a periodically driven, weakly damped FPUT chain. Specifically, near the first mode, we identify a narrow region in the parameter space of driving amplitude and damping where energy is periodically exchanged among a few modes over long timescales. The maximum damping that allows such recurrence decreases rapidly with system size. For example, it is approximately $4.3\times 10^{-3}$ for $N=8$, and about $5\times 10^{-5}$ for $N=32$. These results suggest that in the thermodynamic limit, such recurrence behavior is unlikely to persist. We interpret this phenomenon as a new type of coherent nonlinear dynamics enabled by a delicate balance between weak driving, weak damping, and the system’s modal structure.
We would like to clarify the use of the term recurrence. Strictly speaking, in a driven-dissipative system, the long-period behavior observed in the modal energies should not be interpreted as recurrence in the Hamiltonian sense–that is, a return to the initial state, since the system no longer evolves on a constant-energy manifold. The long-time dynamics of driven-dissipative systems is governed by attractors, and the influence of the initial condition decays during the transient stage. Accordingly, the phenomenon reported in this work is more precisely described as a long-period energy modulation or a coherent energy exchange cycle, rather than a genuine recurrence. Nevertheless, in order to emphasize its phenomenological connection with the original FPUT recurrence, we retain the term recurrence throughout the paper, while acknowledging this distinction.
The paper is organized as follows. Section II introduces the model and the methodology used in our numerical simulations. Section III presents the numerical results, including representative recurrence phenomena and the dependence of the system’s response on driving frequency, driving amplitude, damping, and chain length. Finally, in Sec. IV, we first briefly discuss our numerical simulations from the perspective of soliton dynamics in the KdV equation and in the Toda lattice. We then examine the connection between the observed recurrence behavior and the recently explored concept of time crystals. Appendices A and B provide supplementary materials on recurrence under various driving conditions in the $\alpha$-FPUT chains and on recurrence in the $\beta$-FPUT chains, respectively.
@@ II Model and methodology for our numerical simulations
[…]
@@
We integrate Eq. (1) using a Runge-Kutta algorithm (verified with both MATLAB’s ode45 and Python’s scipy.integrate.solve_ivp, yielding consistent results), with initial conditions $q_{n}=0$ and $\dot{q}_{n}=0$. In the original FPUT problem, Runge-Kutta methods are generally not used because they are not symplectic and therefore do not guarantee long-term energy conservation. In our driven–damped model, however, symplectic algorithm is not required. To ensure numerical accuracy, the solver’s relative and absolute error tolerances were reduced to $10^{-8}$ and $10^{-12}$, respectively, which are significantly smaller than the default settings.
we have monitored the hamiltonian of the undriven, undamped $\alpha$-FPUT as the “total energy”
@@
[…]
Figure 2: Representative recurrence phenomenon. (a–c) Simulation results for the $N=8$ chain with $\eta=2.5\times 10^{-3}$, driving amplitude $F=5\times 10^{-3}$, and driving frequency $\omega=0.347$ (close to the first mode $\omega_{1}=0.3473$). The horizontal axis represents time in units of the driving period $T_{0}$. Panel (c) shows the evolution of modal energies over a time span of $5\times 10^{3}$ driving periods. Due to the high density of curves, the traces appear as a filled region; this representation is intended solely to demonstrate the long-time persistence of the recurrence phenomenon. Local details are successively magnified in panels (b) and (a). (d) Fast Fourier transform (FFT) of the steady-state response shown in panel (c).
@@ III Numerical results
Figure 3: (a) Total energy spectra of the $N=8$ chain as a function of the driving frequency near the first mode $\omega_{1}=0.3473$, $\eta=2.5\times 10^{-3}$, for different driving amplitudes. (b) The same data for $F=5\times 10^{-3}$ shown separately, with a linear scale on the vertical axis. (c) Enlarged view of the frequency range where the recurrence phenomena occur, obtained from a finer frequency scan of the section shown in (b). (d–f) Time-domain simulations for three representative driving frequencies, sampled once per driving period $T_{0}$. (g–i) Enlarged views of the shaded regions in panels (d–f), sampled at intervals of $0.05\,T_{0}$. In panels (g–i), the energies of all modes exhibit oscillations with a period of $0.5\,T_{0}$, indicating that all modes oscillate at the driving frequency.
@@ III.1.2 Energy spectra versus driving frequency
Figure 3(a) shows the total energy response as a function of the driving frequency near the first mode $\omega_{1}=0.3473$, for different driving amplitudes. Specifically, for each set of control parameters, Eq. 1 was integrated for $N$ driving periods $T_{0}$ (typically $N=2000$) until transients had decayed. The total energy was then stroboscopically sampled with period $T_{0}$ over the following $500\,T_{0}$, and these data points are plotted in Fig. 3(a). To maintain consistency, the first sampled point for each frequency scan was taken at the maximum total energy within one $T_{0}$ period.
As the driving amplitude increases, nonlinear effects become more pronounced, giving rise to secondary resonances near the primary peaks, accompanied by jump phenomena in their vicinity. The recurrence phenomena occur in the vicinity of these jumps, which is more clearly seen in Fig. 3(b) and (c). On both sides of the recurrence regions, the system still exhibits normal steady-state oscillations synchronized with the driving period (Figs. 3(d) and (f)). It is worth noting that, regardless of whether recurrence occurs, the system still oscillates at the driving frequency. In particular, every individual mode of the system oscillates at the driving frequency, as shown in Fig. 3(g–i).
@@
[…]
@@ III.1.3 Energy spectra versus driving amplitude
Figure 7: Recurrence regions in the driving amplitude–damping plane for chains of length $N=8$, $16$, and $32$, each driven at a frequency close to its first normal mode. Red filled circles and black filled squares indicate the boundary points of each region, and the connecting lines are drawn only to guide the eye. Note the logarithmic scale: the parameter space in which recurrence is observed becomes significantly narrower for $N=32$ compared to $N=8$. The recurrence regions have an apparent lower boundary at $\eta=10^{-5}$ due to the artificial limit on the damping; smaller values would require much longer simulations to reach a steady state.
@@ III.2 Recurrence regions versus chain Length
IV Discussion and Conclusion
We report numerical evidence of recurrence phenomena in driven–damped $\alpha$-FPUT chains. We also performed similar investigations for driven–damped $\beta$-FPUT chains. For a chain of length $N=32$ with $\beta=8$, analogous recurrence was observed (see Appendix B for details). The choice of $N=32$ and $\beta=8$ is consistent with the parameter settings used in the original work of Fermi et al. [1].
Moreover, for the undriven and undamped FPUT system, recurrence has been famously related to soliton dynamics in the KdV equation or in the Toda lattice[26, 27, 28]. Both the KdV equation and the Toda lattice are integrable models that admit analytical descriptions. In contrast, the system studied here is a finite-length FPUT chain subject to both external driving and damping, a setting that fundamentally breaks integrability. In particular, the reduction of the FPUT lattice to the KdV equation is only valid under a continuum approximation with specific weakly nonlinear, long-wavelength, and unidirectional asymptotic scalings. As a result, the soliton-based interpretation of recurrence may not directly applicable to the present system. From a theoretical perspective, the most feasible analytical route would be a perturbative treatmen[29]. However, even for a short chain with $N=8$, such an approach becomes mathematically cumbersome and technically demanding due to the presence of both driving and damping. Consequently, deriving analytical approximations that describe the system’s behavior inside the recurrence regions remains challenging, providing a natural direction for future theoretical exploration.
@@ Acknowledgments
This research was supported by the Research Project Supported by Shanxi Scholarship Council of China (Grant No. 2023-033), and the Fundamental Research Program of Shanxi Province (Grant No. 202303021221071).
Appendix A Recurrence under Various Driving Conditions
Figure 8 shows the recurrence behavior of an $N=8$ FPUT chain under uniform, single-site, and mode 1-matched driving at $\omega=0.347$. Figure 9 presents results for FPUT chains under a staggered driving, $F_{n}=a\,(-1)^{n}\cos(\Omega t)$, with the same driving amplitude $F=0.15$ and damping $\eta=2.5\times 10^{-3}$. For staggered driving, the site-dependent driving amplitude distribution exhibits different symmetries with respect to the chain center, depending on the chain length: it is even-symmetric for odd $N$ and odd-symmetric for even $N$ (Fig. 9(e)). Consequently, when driving near the frequency of the first normal mode, which has even symmetry, direct excitation of mode 1 is possible only for chains with odd length $N$ (see Fig. 9(c–d), where the nonlinear coupling coefficient is set to zero, $\alpha=0$). Figures 9(a) and 9(b) correspond to driving near the first normal mode for $N=9$ and $N=8$, respectively. A recurrence phenomenon similar to that observed under uniform driving is clearly present for the odd-length chain, whereas no such recurrence is observed for the even-length chain. In fact, staggered driving is highly inefficient for chains with even $N$. As shown in Fig. 9(b), the steady-state total energy reaches only about $0.027$, which is much smaller than that in Fig. 9(a) and is almost indistinguishable from the linear-chain response in Fig. 9(d). This indicates that the system remains in a weakly excited regime where nonlinear coupling plays a negligible role.
Figure 8: Recurrence behavior of an $N=8$ FPUT chain under uniform, single-site, and mode 1-matched driving at $\Omega=0.347$, with damping $\eta=2.5\times 10^{-3}$ and nonlinear coupling coefficient $\alpha=0.25$.
Figure 9: (a) and (b) Staggered driving, $F_{n}=a\,(-1)^{n}\cos(\Omega t)$, for $N=9$ (odd, recurrence) and $N=8$ (even, no recurrence), driven near the first normal mode, $\Omega=0.347$ and $0.313$, respectively, with $\eta=2.5\times 10^{-3}$ and $\alpha=0.25$. (c) and (d) show the excitation behavior of the linear chain ($\alpha=0$): mode 1 dominates for $N=9$, while odd modes are forbidden for $N=8$. The insets in panels (b–d) display the energy distribution among the normal modes at time $t=500\mathrm{T_{0}}$. Note that mode 8 eventually dominates for $N=8$ because it is the best spatial match to the staggered driving. (e) The spatial symmetry of the driving amplitude distribution and the normal modes.
Appendix B Recurrence in the driven damped $\beta$-FPUT chains
For a chain of length $N=32$ with $\beta=8$, we performed scans of the total-energy spectra as functions of the driving frequency in the vicinity of the first normal-mode frequency ($\omega_{1}=0.09516$). The scans were carried out under both uniform driving (symmetric under lattice reflection, where excitation of even modes is forbidden) and single-site driving (non-symmetric under lattice reflection, where excitation of even modes is allowed). Within the explored parameter ranges, analogous recurrence behavior was observed (see Figs. 10 and 11 for details).
It can be seen that Figs. 10 and 11 are nearly identical, despite corresponding to different driving parameters. This indicates that the recurrence phenomenon is robust to different forms of the driving term. When the system reaches comparable total energies, the recurrence patterns under different driving forms are nearly the same.
Furthermore, in both driving protocols of the $\beta$-FPUT chain, we find that the system’s energy is primarily concentrated in the four odd modes 1, 3, 5, and 7. Under uniform driving, the energies of the even modes remain strictly zero. Under single-site driving, the even modes are excited, but their energies are significantly smaller in magnitude compared with the four dominant odd modes.
To illustrate this point, in the normal-mode coordinates, the equations of motion can be written as follows:
$\ddot{Q}_{k}=-\omega_{k}{}^{2}Q_{k}-\frac{\beta}{4}\sum_{i,j,m=1}^{N}B_{k,ijm}Q_{i}Q_{j}Q_{m}+f_{k}(\Omega t)+\sum_{l}^{N}\eta_{l}\dot{Q}_{l}$ (6)
where the term $f_{k}(\Omega t)$ corresponds to the external driving, $\eta_{l}\dot{Q}_{l}$ represents dissipation, and the mode-coupling coefficients are given by:
$B_{k,ijm}=\frac{\omega_{k}\omega_{i}\omega_{j}\omega_{m}}{2(N+1)}\sum_{\pm}\left[\delta_{k,\pm i\pm j\pm m}-\delta_{k\pm i\pm j\pm m,\pm 2(N+1)}\right]$ (7)
Since the periodic driving frequency is close to the first normal-mode frequency, if mode coupling is neglected, the energy injected directly by the external driving is concentrated primarily in mode 1, while direct excitation of the other modes diminishes with increasing frequency detuning. Analogous to the original (undriven) FPUT problem, this situation is, to some extent, equivalent to having the initial energy highly concentrated in mode 1, that is, the initial values satisfy $Q_{k\neq 1}\ll Q_{1}$ and $\dot{Q}_{k}=0$. According to Eq. 7, a necessary condition for the mode-coupling coefficient $B_{\text{even},ijk}$ to be nonzero is that at least one of $i$, $j$, or $k$ is even. Therefore, the corresponding coupling term
$B_{\text{even},ijm}Q_{i}Q_{j}Q_{m}$
remains small, since at least one of the coordinates $Q_{i}$, $Q_{j}$, or $Q_{m}$ is small. As a consequence, the coordinate amplitudes of the even modes remain negligibly small throughout the evolution.
In contrast, the first few odd modes can gradually acquire a finite fraction of the total energy through the coupling terms
$B_{3,111}\,Q_{1}Q_{1}Q_{1},\quad B_{5,113}\,Q_{1}Q_{1}Q_{3},\quad\text{and}\quad B_{7,115}\,Q_{1}Q_{1}Q_{5},$
allowing these modes to progressively share energy during the system’s evolution.
Figure 10: Uniform driving ($F_{n}=F\cos(\Omega t)$) with $N=32$, $\beta=8$, and $\eta=10^{-4}$. (a) Total-energy spectra as a function of the driving frequency near the first normal-mode frequency, $\omega_{1}=0.09516$, for different driving amplitudes. (b) Time-domain simulations corresponding to the multivalued total energy in panel (a) ($\Omega=0.099$, $F=2.5~\times 10^{-4}$), illustrating a chaotic steady state. (c) Enlarged view of panel (a) showing the local structure. (d) Total-energy spectra as a function of the driving amplitude at fixed driving frequency $\Omega=0.0955$. (e–f) Recurrence phenomena observed in the two multivalued regions in panel (d). In the insets, the vertical axis is plotted on a logarithmic scale to facilitate comparison of the modal energy magnitudes.
Figure 11: Single site driving ($F_{n}=F\,\delta_{n,4}\cos(\Omega t)$) with $N=32$, $\beta=8$, and $\eta=10^{-4}$. (a) Total-energy spectra as a function of the driving frequency near the first normal-mode frequency, $\omega_{1}=0.09516$, for different driving amplitudes. (b) Time-domain simulations corresponding to the multivalued total energy in panel (a) ($\Omega=0.0975$, $F=10~\times 10^{-3}$), illustrating a chaotic steady state. (c) Enlarged view of panel (a) showing the local structure. (d) Total-energy spectra as a function of the driving amplitude at fixed driving frequency $\Omega=0.0955$. (e–f) Recurrence phenomena observed in the two multivalued regions in panel (d). In the insets, the vertical axis is plotted on a logarithmic scale to facilitate comparison of the modal energy magnitudes.